# adF: a Novel System for Measuring Web Fingerprinting through Ads

### v2

@@ 2.3. Popular browser’ approaches to counter fingerprinting
- Brave: While Brave is subject to web fingerprinting, it is not exposed to web fingerprinting from ads since the advertisements shown by brave are simple text notifications where code cannot be embedded. To mitigate web fingerprinting, Brave blocks several APIs (such as Canvas, Battery Status, WebGL, and Web Bluetooth, among others). These functions are natively integrated into the browser (brave_fp_protection_mode, ).
- Tor: It is the browser proposing the most robust solution against fingerprinting. Its approach is to build the same fingerprint for all devices in the web (laperdrix2019torbrowser, ). However, this approach requires blocking some essential functions, which affects the user’s browsing experience (khattak2016you, ).
- Safari: In 2018, Apple released a version of Safari with an anti-fingerprinting setting already built into the browser and thus requiring no user action (cameron_2018_appleagainstfp, ). In particular, Safari does not support specific APIs such as Web Bluetooth, Device Memory, or User Permission State (e.g., camera, geolocation, or microphone). It also removes detailed information on the user agent or on the type of graphic card used.
- Firefox: Mozilla Firefox offers user-configurable privacy options through fingerprint controls. In 2019, Firefox enabled the Enhanced Tracking Protection feature by default as part of its settings (enhanced_tracking_protection_mozillawiki, ). Today, this functionality offers opt-in fingerprint protection preventing the collection of information related to time zone, installed fonts, window size preferences, and operating system versions (among others).
- Chrome: It is the browser with the largest market share. Google announced in August 2019 that they would introduce fingerprint-blocking features in Chrome but without providing an expected release date (schuh_2019_chrome_privacy, ). It was not until early 2020 that Google revealed their plans: reduce the granularity of the information provided by the user agent, which is sent through every HTTP request by default (user_agent_reduction_privacysandbox, ). This functionality is still in the experimental phase.
@@
Finally, in the case of mobile apps, operating systems such as iOS and Android are implementing significant changes to increase privacy and security by restricting access to data that could be used for fingerprinting (jones2023, ; macrumors2023, ; yodelmobile2023, ). Based on our results, we find that iOS and Android restrict access to information that might be used for fingerprinting from their mobile applications.
3. Threat Model for the adF
In this section, we detail the specific threat model for adF, the system we present in Section 4 to evaluate the vulnerability of different device configurations to web fingerprinting.
3.1. System model
@@
3.2. Attacker model
We consider a remote attacker who wants to track a web user through web fingerprinting techniques embedded in advertisements. In the threat model, the attacker is a malicious ad-tech provider or any other malicious player with access to advertising platforms capable of running ad campaigns. The attacker can add a JS pixel into ads, which is widely accepted and not considered an illegal technique in ad-tech. This makes it relatively easy for the attacker to perform the attack without the need for third-party collaboration or black-hat hacking techniques to gain access to any third-party venue (e.g., a web-page). This approach significantly lowers the barriers to conducting such attacks, allowing potential reach to tens of millions of users for fingerprinting purposes.
In contrast, the attacker model defined in previous works in the literature (alaca2016device, ; eckersley2010, ; boda2011user, ; mowery2012pixel, ; fifield2015fingerprinting, ; laperdrix2016beauty, ; gomezBoix2018hiding, ; laperdrix2021fingerprinting, ; vastel2018fp, ; schwarz_2019, ; rizzo2021unveiling, ; Konstantinos2022Chronos, ; andriamilanto2021largescale, ) requires the attacker to gain access to one or multiple web-pages to inject a fingerprinting script. This requires either the cooperation of a third party, such as the web-page owner—who may be hesitant due to data protection concerns (e.g., GDPR compliance)—or illegally gaining access to the website by black-hat hacking techniques. To reach tens of millions of users, the accessed websites should be relatively popular, which (in principle) makes it even harder to reach an agreement with their owners due to potential bad reputation implications or hack them (since it is assumed popular web-pages have better security measures in place compared to unpopular ones). Therefore, traditional threat models presented in the literature depict a much higher barrier to performing a fingerprinting attack compared to the alternative using ads discussed in this paper.
4. adF: A system to measure web fingerprinting from ads
@@ 4.1. Data Collector
To collect attributes from the device, we develop an adTag, a JavaScript (JS) code, which can be embedded into any ad format capable of running a script. Our adTag is executed upon the rendering process of the ad on a device. The use of adTags is a widespread practice in online advertising. Moreover, the research community has also used adTags to analyze transparency aspects of online advertising (callejo2019q, ; 8932475, ) and to conduct network measurements (callejo2017opportunities, ; callejo2019measuring, ).
@@
Our adTag captures a comprehensive set of up to 191 diverse attributes from the device, OS, and browser where the ad is displayed. Specifically, it leverages JavaScript APIs within browsers to retrieve information. While some APIs necessitate permissions for access, such as the microphone (get_microphone_permission, ), camera (mediaDevices_getUserMedia, ), or location (geolocation_API, ), the majority of them are openly accessible through our adTag. The attributes retrieved by our adTag via these APIs are stored on a server within our domain, where we process them and construct the fingerprint using the Fingerprint Constructor. Table 4 in Appendix A provides the complete list of attributes considered, along with their respective sources.
@@
An initial analysis of the 191 considered attributes, based on tens of thousands of collected samples, reveals that some of these attributes are not suitable for creating unique fingerprints. Therefore, our Data Collector does not retrieve attributes that meet any of the following criteria: 1) the attribute reports a constant value (e.g., Bars settings: menubar); 2) the value of the attribute is unstable over time (e.g., Network Information: downlink); 3) the value is unreliable (e.g., Window navigator.onLine); 4) most popular browsers do not support the attribute (e.g., WebHID API). After applying this filtering, the final version of the Data Collector retrieves a total of 66 attributes from the initial 191. These attributes are highlighted in bold in Table 4. Note that some of these 66 attributes are joined to form meta-attributes defined as a concatenation of several individual attributes. Table 4 also details the mapping of attributes to meta-attributes.
@@
In the case of mobile apps, just a subset of 35 of these 66 attributes are considered and thus used in our analyses. These are identified in Table 4 with an asterisk (*).
@@
Ads are expected to be exposed on a device for a limited time. Usually, a few seconds (exposure_time_ads, ). Therefore, our adTag must collect the 66 (35) attributes and deliver them to the backend server before the user closes the website (or the mobile app) where the ad is displayed. To this end, we use AJAX (Asynchronous JavaScript and XML) (ajax_developerguides_mdn_2022, ), which allows the asynchronous execution of code so that the most time-consuming operations do not block the rest of the script’s tasks. By doing so, we ensure that for more than 95% of the ads we received a response from the browser (mobile app) including the complete list of attributes. Note that we discard incomplete fingerprint samples to avoid generating fingerprints that may differ due to these uncollected attributes.
Figure 1. adF - our web fingerprinting system.
4.2. Fingerprint Constructor
@@
This module aims to build a device’s fingerprint using the 66 (35) attributes collected by the Data Collector from the browser (mobile app). To this end, the Fingerprint Constructor first generates a string concatenating the value of these 66 (35) attributes. On average, this string has a size of 657 Bytes among the samples available in our dataset (as described in Section 5). We could use this string as a fingerprint of the device. However, for enhanced computational efficiency, we apply an MD5 hash function to create an alphanumeric fingerprint consisting of 512 bits for each device. The use of a hash function ensures a very low collision probability, and the resulting fingerprint’s reduced size enables rapid computational processing.
4.3. Fingerprint Classifier
@@
The fingerprint delivered by the Fingerprint Constructor is actionable, for instance, to track a device in the web, only if it is unique. Consequently, the final component of our system is a binary (fingerprint) classifier. Its output, i.e., the dependent variable, is a binary indicator that determines whether a fingerprint is unique (1) or non-unique (0). It receives as input (i.e., independent variables) the 66 (35) attributes from the Data Collector for browser (mobile app) samples.
@@
The Fingerprint Classifier, like any classifier, requires training. To achieve this, we use the fingerprint datasets, as introduced in Section 5, comprising 161k and 52k real fingerprints from browsers and mobile apps, respectively. Each fingerprint in these datasets is labeled as either unique or non-unique.
The algorithm employed to build the model for our Fingerprint Classifier is XGBoost (eXtreme Gradient Boosting) (xgboost_wiki, ). It has demonstrated the highest performance in our evaluation using a k-fold cross-validation approach when compared to alternative methods, including Logistic Regression, Support Vector Machine, Decision Tree, Random Forest, Gradient Boosting Machine, as well as various third-party gradient boosting algorithms like LightGBM and CatBoost.
Two essential tasks related to data handling are integral to the design of our classifier. Firstly, we employ the k-nearest neighbor (KNN) imputation technique to address the issue of missing values that are not reported by the browser (or mobile app). KNN is a widely used and effective technique (aydin2021performance, ; scikit_2022, ). Secondly, XGBoost, like many other machine learning algorithms, was primarily designed to work with numerical data [^2] Note that the latest versions of XGBoost offer experimental support for categorical variables. However, this support is still limited (categoricaldata_xgboost_1_7_1, ).. To transform categorical variables into numerical ones, we adhere to best practices recommended by the research community. We utilize one-hot encoding (sklearn_onehotencoder, ) for categorical variables with low cardinality and frequency encoding (countfrequencyencoder_1_0_2, ) for those with high cardinality.
Table 1. Details of the ad campaigns used to populate our fingerprint datasets.
@@
Finally, we choose to develop two separate models, one for browsers and another for mobile apps. This decision is driven by the different number of attributes available in each case, 66 vs. 35. We also evaluate the option of building different independent models for individual device configurations in our datasets (e.g., a specific model for desktop devices running a Windows OS or for mobile Android apps), but we observe that the performance of adF is similar for the general and configuration-specific models, and thus we decide to implement the two referred general models.
5. Measurements and Datasets
To assess the vulnerability of different device configurations within browser-based environments to web fingerprinting, we deploy the adF system, as described earlier, in live ad campaigns. These campaigns are configured within the Sonata Platform, a Demand Side Platform (DSP) operated by TAPTAP Digital (taptap_digital, ). Sonata is a mid-sized DSP that serves tens of millions of daily ads in 15 countries across Europe, Africa, and the Americas. The Data Collector retrieves the 66 attributes (highlighted in Table 4) from each of the ads delivered through these ad campaigns. Each sample is then fed into the Fingerprint Constructor and the Fingerprint Classifier to obtain the device’s fingerprint and determine whether it is unique or not. In addition, when available, we also collect the device’s advertising ID provided by the DSP as a feature.
@@ 5. Measurements and Datasets
We collect a second dataset, comprising 975k samples from mobile apps, distinct from the browser-based data. This dataset is referred to as the raw-apps dataset. The samples in this dataset contain the same fields as previously described, with specific differences: (1) the field device’s type is always mobile in this case; (2) the field device’s OS can take only two values, iOS or Android; (3) the field browser identifies, in this case, the WebView of the Android or iOS system; (4) the data structure attributes includes information for 35 attributes.
@@
Finally, it is well-known that the market share of OSes and browsers is imbalanced. On the one hand, Chrome holds a two-thirds share of browser installations on both desktop and mobile devices. On the other hand, Windows OS dominates the desktop market with a share over 72,1% while the mobile device market is primarily split between Android (71%) and iOS (28,3%). To ensure that our dataset adequately represents the most popular browsers and OSes, we utilize the targeting capabilities provided by Sonata DSP to conduct specific ad campaigns tailored to particular OSes and browsers. Additionally, we run specific campaigns for mobile apps. The details of each campaign can be found in Table 1. In addition, Figure 2 shows the number of samples in our fingerprint-browsers dataset across device configurations accounting for at least 725 samples. For statistical representativeness reasons, these are the configurations we consider in our analyses in the rest of the paper. In the case of mobile apps, our fingerprint-apps dataset includes 20,66% (10686) and 79,34% (41031) samples from iOS and Android apps, respectively.
Figure 2. x-axis: $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ performance of adF system for all the considered browser-based device and mobile app configurations; y-axis: number of samples for the browser and mobile app configurations considered for the analyses conducted in the paper.
6. Vulnerability and Accuracy metrics

### v3

@@ II-C Popular browser’ approaches to counter fingerprinting
- Brave: While Brave is subject to web fingerprinting, it is not exposed to web fingerprinting from ads since the advertisements shown by brave are simple text notifications where code cannot be embedded. To mitigate web fingerprinting, Brave blocks several APIs (such as Canvas, Battery Status, WebGL, and Web Bluetooth, among others). These functions are natively integrated into the browser [brave_fp_protection_mode].
- Tor: It is the browser proposing the most robust solution against fingerprinting. Its approach is to build the same fingerprint for all devices on the web [laperdrix2020fpSurvey]. However, this approach requires blocking some essential functions, which affects the user’s browsing experience [khattak2016you].
- Safari: In 2018, Apple released a version of Safari with an anti-fingerprinting setting already built into the browser and thus requiring no user action [cameron_2018_appleagainstfp]. In particular, Safari does not support specific APIs such as Web Bluetooth, Device Memory, or User Permission State (e.g., camera, geolocation, or microphone). It also removes detailed information on the user agent or on the type of graphic card used.
- Firefox: Mozilla Firefox offers user-configurable privacy options through fingerprint controls. In 2019, Firefox enabled the Enhanced Tracking Protection feature by default as part of its settings [enhanced_tracking_protection_mozillawiki]. Today, this functionality offers opt-in fingerprint protection preventing the collection of information related to time zone, installed fonts, window size preferences, and operating system versions (among others).
- Chrome: It is the browser with the largest market share. Google announced in August 2019 that they would introduce fingerprint-blocking features in Chrome but without providing an expected release date. It was not until early 2020 that Google revealed their plans: reduce the granularity of the information provided by the user agent, which is sent through every HTTP request by default [user_agent_reduction_privacysandbox]. This functionality is still in the experimental phase.
@@
Finally, in the case of mobile apps, operating systems such as iOS and Android are implementing significant changes to increase privacy and security by restricting access to data that could be used for fingerprinting [yodelmobile2023]. Based on our results, we find that iOS and Android restrict access to information that might be used for fingerprinting from their mobile applications.
III Threat Model for the adF
In this section, we detail the specific threat model for adF, the system we present in Section IV to evaluate the vulnerability of different device configurations to web fingerprinting.
III-A System model
@@
III-B Attacker model
We consider a remote attacker who wants to track a web user through web fingerprinting techniques embedded in advertisements. In the threat model, the attacker is a malicious ad-tech provider or any other malicious player with access to advertising platforms capable of running ad campaigns. The attacker can add a JavaScript (JS) pixel into ads, which is widely accepted and not considered an illegal technique in ad-tech. This makes it relatively easy for the attacker to perform the attack without the need for third-party collaboration or black-hat hacking techniques to gain access to any third-party venue (e.g., a web-page). This approach significantly lowers the barriers to conducting such attacks, allowing potential reach to tens of millions of users for fingerprinting purposes.
In contrast, the attacker model defined in previous works in the literature [eckersley2010, boda2011user, laperdrix2016beauty, gomezBoix2018hiding, laperdrix2021fingerprinting, andriamilanto2021largescale] requires the attacker to gain access to one or multiple web-pages to inject a fingerprinting script. This requires either the cooperation of a third party, such as the web-page owner—who may be hesitant due to data protection concerns (e.g., GDPR compliance)—or illegally gaining access to the website by black-hat hacking techniques. To reach tens of millions of users, the accessed websites should be relatively popular, which (in principle) makes it even harder to reach an agreement with their owners due to potential bad reputation implications or hack them (since it is assumed popular web-pages have better security measures in place compared to unpopular ones). Therefore, traditional threat models presented in the literature depict a much higher barrier to performing a fingerprinting attack compared to the alternative using ads discussed in this paper.
IV adF: A system to measure web fingerprinting from ads
@@ IV-A Data Collector
To collect attributes from the device, we develop an adTag, a JavaScript (JS) code, which can be embedded into any ad format capable of running a script. Our adTag is executed upon the rendering process of the ad on a device. The use of adTags is a widespread practice in online advertising. Moreover, the research community has also used adTags to analyze transparency aspects of online advertising [callejo2019q] and to conduct network measurements [callejo2017opportunities].
@@
Our adTag captures a comprehensive set of up to 191 diverse attributes from the device, OS, and browser where the ad is displayed. Specifically, it leverages JavaScript APIs within browsers to retrieve information. While some APIs necessitate permissions for access, such as the microphone, camera, or location, the majority of them are openly accessible through our adTag. The attributes retrieved by our adTag via these APIs are stored on a server within our domain, where we process them and construct the fingerprint using the Fingerprint Constructor. Table IV in Appendix A provides the complete list of attributes considered, along with their respective sources.
Figure 1: adF visual representation, our web fingerprinting system.
@@
An initial analysis of the 191 considered attributes, based on tens of thousands of collected samples, reveals that some of these attributes are not suitable for creating unique fingerprints. Therefore, our Data Collector does not retrieve attributes that meet any of the following criteria: 1) the attribute reports a constant value (e.g., Bars settings: menubar); 2) the value of the attribute is unstable over time (e.g., Network Information: downlink); 3) the value is unreliable (e.g., Window navigator.onLine); 4) most popular browsers do not support the attribute (e.g., WebHID API). After applying this filtering, the final version of the Data Collector retrieves a total of 66 attributes from the initial 191. These attributes are highlighted in bold in Table IV. Note that some of these 66 attributes are joined to form meta-attributes defined as a concatenation of several individual attributes. Table IV also details the mapping of attributes to meta-attributes.
@@
In the case of mobile apps, just a subset of 35 of these 66 attributes are considered and thus used in our analyses. These are identified in Table IV with an asterisk (*).
@@
Ads are expected to be exposed on a device for a limited time. Usually, a few seconds [exposure_time_ads]. Therefore, our adTag must collect the 66 (35) attributes and deliver them to the backend server before the user closes the website (or the mobile app) where the ad is displayed. To this end, we use AJAX (Asynchronous JavaScript and XML), which allows the asynchronous execution of code so that the most time-consuming operations do not block the rest of the script’s tasks. By doing so, we ensure that for more than 95% of the ads we received a response from the browser (mobile app) including the complete list of attributes. Note that we discard incomplete fingerprint samples to avoid generating fingerprints that may differ due to these uncollected attributes.
IV-B Fingerprint Constructor
@@
This module aims to build a device’s fingerprint using the 66 (35) attributes collected by the Data Collector from the browser (mobile app). To this end, the Fingerprint Constructor first generates a string concatenating the value of these 66 (35) attributes. On average, this string has a size of 657 (489) Bytes among the samples available in our dataset (as described in Section V). We could use this string as a fingerprint of the device. However, for enhanced computational efficiency and integrity, we apply the SHA-256 cryptographic hash function. This creates a 256-bit alphanumeric fingerprint for each device. The use of a hash function ensures a negligible collision probability, and the resulting fingerprint’s fixed, reduced size enables rapid computational processing.
IV-C Fingerprint Classifier
@@
The fingerprint delivered by the Fingerprint Constructor is actionable, for instance, to track a device on the web, only if it is unique. Consequently, the final component of our system is a binary (fingerprint) classifier. Its output, i.e., the dependent variable, is a binary indicator that determines whether a fingerprint is unique (1) or non-unique (0). It receives as input (i.e., independent variables) the 66 (35) attributes from the Data Collector for browser (mobile app) samples.
@@
The Fingerprint Classifier, like any classifier, requires training. To achieve this, we use the fingerprint datasets, as introduced in Section V, comprising 161k and 52k real fingerprints from browsers and mobile apps, respectively. Each fingerprint in these datasets is labelled as either unique or non-unique.
The algorithm we employ to build the model for our Fingerprint Classifier is XGBoost (eXtreme Gradient Boosting), which outperformed alternative methods-such as Logistic Regression, Support Vector Machine, Decision Tree, Random Forest, Gradient Boosting Machine, as well as third-party gradient boosting algorithms like LightGBM and CatBoost-in our evaluations. To robustly estimate model performance while maximizing the use of our labelled data (samples with available advertising ID, see Section V), we adopt a 10-fold stratified cross-validation procedure-an approach to offer an effective balance between bias and variance for moderate-size datasets [arlot2010survey]. Specifically, the dataset is randomly partitioned into 10 equal-sized folds, where in each iteration, the model is trained on 9 folds and tested on the remaining fold. This approach ensures that every fingerprint is used once for testing without any overlap between training and validation, thereby preventing potential data leakage. Moreover, to prevent correlation leakage at the device level, the 10-fold split is also grouped by advertising ID: all samples linked to the same Ad-ID are assigned to the same fold. The Ad-ID is used solely for ground-truth construction and fold assignment and is never included as a model input.
Beyond standard i.i.d. cross-validation, we also assess out-of-distribution (OOD) generalization by (i) leaving one {browser, OS} configuration out (LOCO) and (ii) temporal hold-outs across campaigns, always grouping folds and splits by Ad-ID to prevent device-level leakage. Detailed protocols and results are reported in Appendix B.
Two essential tasks related to data handling are integral to the design of our classifier. Firstly, we apply the k-nearest neighbor (KNN) imputation technique to address structurally missing values—i.e., attributes that are consistently unreported by certain browser or mobile app configurations due to the inherent behaviour or policy of those environments, not accidental failures. Only numerical attributes are imputed, either originally continuous or transformed from high-cardinality categorical variables via frequency encoding. We use $k=5$ and the nan_euclidean distance metric, which computes distances over the available attributes only, making it well suited for partially missing fingerprints. Secondly, since XGBoost is primarily designed to work with numerical inputs, [^2] Note that the latest versions of XGBoost offer experimental support for categorical variables. However, this support is still limited. we transform categorical variables following best practices: one-hot encoding for low-cardinality attributes and frequency encoding for high-cardinality ones. Frequency-encoded attributes yield continuous numerical values that reflect the statistical distribution of categories in the dataset, making them suitable for distance-based imputation methods.
Table I: Details of the ad campaigns used to populate our fingerprint datasets.
@@
Finally, we choose to develop two separate models, one for browsers and another for mobile apps. This decision is driven by the different number of attributes available in each case, 66 vs. 35. We also evaluate the option of building different independent models for individual device configurations in our datasets (e.g., a specific model for desktop devices running a Windows OS or for mobile Android apps), but we observe that the performance of adF is similar for the general and configuration-specific models, and thus we decide to implement the two referred general models. To further validate the robustness of these models, we compute additional evaluation metrics. The models achieve ROC AUC values of 90% and 88%, and balanced accuracy scores of 78% and 80% for browser-based and mobile app fingerprints, respectively.
V Measurements and Datasets
To assess the vulnerability of different device configurations within browser-based environments to web fingerprinting, we deploy the adF system, as described earlier, in live ad campaigns. These campaigns are configured within the Sonata Platform, a Demand Side Platform (DSP) operated by TAPTAP Digital. Sonata is a mid-sized DSP that serves tens of millions of daily ads in 15 countries across Europe, Africa, and the Americas. The Data Collector retrieves the 66 attributes (highlighted in Table IV) from each of the ads delivered through these ad campaigns. Each sample is then fed into the Fingerprint Constructor and the Fingerprint Classifier to obtain the device’s fingerprint and determine whether it is unique or not. In addition, when available, we also collect the device’s advertising ID provided by the DSP as a field.
@@ V Measurements and Datasets
We collect a second dataset, comprising 975k samples from mobile apps, distinct from the browser-based data. This dataset is referred to as the raw-apps dataset. The samples contain the same fields as previously described, with specific differences: (1) the field device’s type is always mobile in this case; (2) the field device’s OS can take only two values, iOS or Android; (3) the field browser identifies, in this case, the WebView of the Android or iOS system; (4) the data structure attributes includes information for 35 attributes.
@@
Finally, it is well-known that the market share of OSes and browsers is imbalanced. On the one hand, Chrome holds a two-thirds share of browser installations on both desktop and mobile devices. On the other hand, Windows OS dominates the desktop market with a share over 72.1% while the mobile device market is primarily split between Android (71%) and iOS (28.3%). To ensure that our dataset adequately represents the most popular browsers and OSes, we utilize the targeting capabilities provided by Sonata DSP to conduct specific ad campaigns tailored to particular OSes and browsers. Additionally, we run specific campaigns for mobile apps. The details of each campaign can be found in Table I. In addition, Figure 2 shows the number of samples in our fingerprint-browsers dataset across device configurations accounting for at least 725 samples. For statistical representativeness reasons, these are the configurations we consider in our analyses in the rest of the paper. In the case of mobile apps, our fingerprint-apps dataset includes 20.66% (10686) and 79.34% (41031) samples from iOS and Android apps, respectively.
Figure 2: x-axis: $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ performance of adF system for all the considered browser-based device and mobile app configurations; y-axis: number of samples for the browser and mobile app configurations considered for the analyses conducted in the paper.
VI Vulnerability & Accuracy